OpenStudy (anonymous):

explain in complete sentences how you would use the elimination method to solve the following system of equations and provide the solution to the system . 5x-3y=4 3x-5y=-4

OpenStudy (anonymous):

1)add two equations 2)subtract the two equations then just use the normal method

OpenStudy (anonymous):

this is harder then i thought :P

OpenStudy (anonymous):

To eliminate y first, multiply the first equation by five, and the second by -3. \[5((5x-3y)=4)\] \[-3((3x-5y)=-4)\] you would then get... \[25x-15y=20\] \[9x+15y=12\] subtract, and you get... \[34x=32\] so, x=16/17 to find y, do the same thing, but just cancel the x-terms. OR, substitute x=16/17 in one of the equations. in the end, \[y=2\] hope this helps, let me know if you have any more questions!